OpenStudy (anonymous):

Write an expression to describe the sequence then find the 10th term -3,-6,-9,-12

OpenStudy (anonymous):

-3= -3(1) -6=-3(2) -9=-3(3) -12=-3(4) therefore, the sequence is -3n where n=1,2,3.....

OpenStudy (anonymous):

the 10th term is -3(10) =-30

OpenStudy (anonymous):

Or either an AP..

OpenStudy (anonymous):

a+(n-1)d a=first term d=common difference n=nth term

OpenStudy (lgbasallote):

^identify those @Numbers can you? :)